OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6?

OpenStudy (tukitw):

Since the line is parallel, their gradient are the same. 2x + 3y = 6 To find the gradient, solve for y. 3y = -2x + 6 y = -2x / 3 + 2 By comparing this equation with y = mx + c, where m is the gradient, the gradient for the two lines is -2 / 3. Next, using the point-slope formula, y = m(x - x1) + y1, plug in the corresponding values, with m being the gradient of the line, (x1 , y1) being any coordinate that lies on the line. So, y = (-2 / 3)(x - -2) + 3, which is from the given (-2 , 3), results in y = (-2 / 3)(x + 2) + 3 y = -2x / 3 + 5 / 3. Finally, rearranging the equation, 3y = -2x + 5 2x + 3y = 5 Thus, the equation of the line that passes through (-2 , 3) is 2x + 3y = 5.